OpenStudy (anonymous):

when you have a positive and a negative number for exponts in scientifc notation..do you multiply add or subtract

OpenStudy (kainui):

\[x^3*x^2=(x*x*x)*(x*x)=x*x*x*x*x=x^5\]

OpenStudy (anonymous):

both of those are positive.. my equation i [2.3 times 10^-3] times [3 times 10^8 power